# Correcting Negative Bias in Large Language Models through Negative Attention Score Alignment

### v1

@@ 6.1 Preservation of General Reasoning Abilities
LLaMA3-8B 0.594 / 0.596 0.329 / 0.325 0.235 / 0.227
Mistral-7B 0.415 / 0.408 0.126 / 0.133 0.174 / 0.174
Gemma-7B 0.139 / 0.137 0.257 / 0.251 0.125 / 0.131
Qwen2-7B 0.541 / 0.541 0.478 / 0.481 0.220 / 0.221
In this section, we evaluate the performance of a large language model with NASA applied to general QA reasoning tasks. We assess the general reasoning performance on benchmarks such as MuSiQue, GSM8K, and AR-LSAT using the LLaMA3-8B model both with and without NASA. After verifying whether each model-generated prediction includes an answer in these benchmarks, we utilize GPT-4 to double-check whether the predictions that contain answers indeed correspond semantically to the actual answers. To generate predictions from the model, we utilize the previously mentioned Prompt II (Inquiring Parametric Knowledge). The content of the GPT-4 prompt for prediction verification can be found in Table 13.
As shown in Table 5, the model with NASA applied demonstrates improved results, showing enhanced or maintained general reasoning abilities while effectively tackling negative bias. Since negative bias is a task-specific factor contributing to hallucination in binary decision tasks, its influence on general reasoning QA, like short-answer QA, is relatively low compared to other factors contributing to hallucination such as parametric knowledge. Therefore, improving reasoning ability in binary decision tasks does not necessarily enhance reasoning capabilities in general reasoning QA. This means that addressing bias in binary decision tasks while maintaining performance in general reasoning QA can be considered an advancement in the model’s overall reasoning capability, not overfitted to the binary decision task.
6.2 Generalization to Universal Binary Decision
We analyze the generalization ability for binary decisions by replacing the positive and negative tokens used in our fine-tuning set. Our fine-tuning set consists of “Yes” and “No”, and in this experiment, we test with the tokens “True” and “False”, as well as “Correct” and “Wrong”.
Type Dataset LLaMA3-8B + NASA
True/False StrategyQA 0.855 / 0.828 0.866 / 0.853
MATH 0.559 / 0.452 0.574 / 0.583
AR-LSAT 0.512 / 0.306 0.506 / 0.405
Correct/Wrong StrategyQA 0.852 / 0.838 0.846 / 0.840
MATH 0.545 / 0.660 0.544 / 0.675
AR-LSAT 0.523 / 0.540 0.522 / 0.590
Table 6 presents the results of experiments with LLaMA3-8B on various types of complex reasoning tasks. As shown in the table, ours fine-tuned using NASA achieves semantic-level generalization without overfitting to specific tokens.
6.3 Transferability across Various Instructions
It. Type Dataset LLaMA3-8B + NASA
A MuSiQue 0.686/0.641 0.710/0.700
GSM8K 0.534/0.472 0.543/0.575
AR-LSAT 0.508/0.306 0.507/0.370
B MuSiQue 0.708/0.692 0.736/0.756
GSM8K 0.536/0.611 0.531/0.665
AR-LSAT 0.512/0.382 0.518/0.500
In this work, we maintain a fixed instruction for binary decision-making in yes-no QA throughout the application of our methodology. In this section, we evaluate the robustness of our method when presented with instructions in the inference stage that are different forms of content. We assess the binary decision performance using the LLaMA3-8B model with and without NASA on benchmarks: MuSiQue, GSM8K, and AR-LSAT for paraphrased instruction types A and B. As shown in Table 7, our methodology with NASA applied still demonstrates superior performance compared to the baseline in both prompt types A and B. This supports our claim of instruction robustness for our methodology. Note that datasets in Table 6 and 7 are rephrased version. Fig 7 in Appendix B shows the content of instructions.
6.4 Enhanced Model Calibration
Dataset Mistral-7B + NASA
@@ 6.4 Enhanced Model Calibration
Table 8: Expected calibration error of baseline and ours.
As Fig 2 illustrates, the negative bias of LLMs is also associated with the model’s confidence in its predictions. Research on the calibration between prediction confidence and accuracy is necessary to enhance the trustworthiness of LLMs (Kadavath et al., 2022).
Table 8 shows the expected calibration error (He et al., 2022) when NASA is applied to Mistral-7B. In this paper, we focus on calibration in a complex reasoning yes-no QA dataset. Across five datasets, NASA consistently demonstrates improved calibration. This indicates that NASA effectively mitigates model bias, resulting in better alignment between prediction confidence and accuracy.
@@ 7 Conclusion
We identify a critical issue where large language models exhibit negative bias in binary decision tasks requiring complex reasoning. To address this issue, we propose a negative attention score for formulation and employ it to discover query-agnostic negative heads. By performing parameter-efficient tuning on these heads, we introduce the NASA method, which effectively mitigates the bias problem. Our method not only enhances the performance of the model but also serves as a useful analytical framework from the perspective of interpretability.
@@ Appendix A Data Statistics
Table 9: Data statistics of datasets utilized in our work.
@@
Appendix B Instructions for NASA
@@
Appendix C Details of Binary Decision Data Construction
@@
C.1 GPT-4 Prompts for Negative Example Construction
@@
Appendix D Details of Parametric Sample Selection
As mentioned in Section 4.1, we construct the probing and fine-tuning sets using the HotpotQA dataset, which is a short-answer QA dataset. From the HotpotQA dataset, we use samples that can be correctly answered solely with parametric knowledge, without any context, as our probing samples.
@@ Appendix D Details of Parametric Sample Selection
Appendix E Case Study for Negative Responses
@@
We apply this classification to cases where the model gives a negative response in a binary decision task. Negative responses can be classified as true negative (TN) or false negative (FN), with our focus in this discussion on FN. In this, an FN response classified as high-confident Det-T is the case where the model knows the answer is correct but says "no", while an FN response classified as Non-Det is the case where the model doesn’t know the correct answer and says "no". For high-confident Det-F responses, the model does not indicate unanswerability but rather holds incorrect knowledge. This issue falls outside the scope of our methodology, which focuses on handling negative bias in binary decision tasks rather than general knowledge-intensive tasks.
@@
Although it is clear that the case where the model knows the answer is correct but says "no" is the most problematic, we argue that in a binary decision task, the case where the model doesn’t know the correct answer and says "no" can be related to the issue of negative bias. In a binary decision task where the model must answer "Yes" or "No," instances where the model is uncertain about the correct answer should result in an equal frequency of positive (yes) and negative (no) responses. However, our analysis indicates that existing models tend to output negative responses more confidently and frequently (as shown in Figures 1 and 2 in the paper). This suggests that when the model is uncertain, it predominantly outputs negative responses, which negatively impacts the trustworthiness of these negative responses. Therefore, for models exhibiting such negative bias, Non-Det and low-confident responses should also be partially adjusted.
Table 14 shows the compositions of FN responses from the original models according to the classifications mentioned. It is evident that in all types of reasoning tasks addressed in our work (multi-hop, mathematical, logical reasoning), the FN responses include both case A and B that you mentioned. Specifically, for Gemma, the model often outputs responses indicating that many questions in Multi-hop QA are unanswerable. Please note that these are classified as Det-F in the table.
Appendix F Analysis of Negative Responses Shifting in NASA
@@
In this section, we analyze the ratios of samples which are originally responded to as negative by the original model and subsequently changed to positive by the NASA model. We categorize the ratio cases by each classification group explained in section E.
@@
Table 15 illustrates the ratios of samples shifted from FN to TP. This demonstrates that the Det-T corresponding to the case where the model knows the answer is correct but says "no" generally has the highest rate of change. This finding indicates that our methodology does not indiscriminately add positive bias to the model but rather enhances the model’s binary decision reasoning ability effectively.
@@
Table 16 illustrates the ratios of samples shifted from TN to FP. As discussed in the section on main results, some samples that are incorrectly categorized as TN due to negative bias might shift to FP (Det-F), and this case does not correspond to the model degradation. Actually, among the TN samples, those classified as Deterministic-T with high confidence exhibit a lower shift ratio compared to their Deterministic-F counterparts. This indicates that the model tends to maintain predictions for samples that it can accurately classify. In conclusion, it can be interpreted that NASA has improved reasoning capability in binary decision tasks while maintaining general reasoning capability.
Appendix G Ablation Study
@@
G.1 Freezing Key Projection Weight
@@
The experimental results of NASA and freezing key projection weights are shown in Table 17. In general, both the F1 score and NAS are relatively better for NASA, while the trend in accuracy is not distinctly observable. An exceptional case is Qwen2-7B at AR-LSAT, where NASA’s accuracy and F1 score are significantly lower than those of its counterpart.
G.2 Random Attention Head Tuning
In another ablation study, we experiment by randomly selecting attention heads to be fine-tuned. As shown in Table 18, NAS typically outperforms most others in terms of F1 score and ECE. Notably, during random attention head tuning, model NAS largely diverges in the case of LLaMA3-8B and Mistral-7B. These results imply that carefully selecting attention heads is critical and our framework effectively detects and addresses negatively biased attention heads.
Appendix H Confidence Histogram Analysis
To analyze the confidence of binary decisions between the baseline and our model, we plot confidence histograms for the LLaMA3-8B model, similar to Fig 2. In Fig 5, we observe that NASA influences the prediction confidence of LLMs in binary decision tasks. In the baseline model, except for StrategyQA, we frequently observe the return of negative responses with high confidence. On the other hand, after incorporating NASA, the model demonstrates a reduction in the frequency and confidence gap between positive and negative responses.
@@ Appendix H Confidence Histogram Analysis
Appendix I Samples of Attention Heads
Fig 6 presents the real examples of attention weights of heads in LLaMA3-8B-Instruct model for a GSM8K rephrased sample. Fig 7 presents the real examples of attention weights of heads in LLaMA3-8B-Instruct model with NASA for a same GSM8K rephrased sample.
Appendix J Computation
@@ Appendix I Samples of Attention Heads
Appendix K Licenses
StrategyQA, MuSiQue, MetaMATH, and AR-LSAT are under the license of MIT license, CC-BY-4.0 license, MIT License, and MIT license, respectively. LLaMA 3, Mistral, Gemma, Qwen, and GPT-4 are under the license of META LLAMA 3 COMMUNITY LICENSE AGREEMENT, Apache License 2.0, Gemma, Apache License 2.0, and OpenAI, respectively.
Appendix L Usage of AI Writing Assistance
@@
Table 10: Instructions used NASA process in our work and Table 7.
@@
Convert the given question to the binary Yes-No question based on the context about the given question and the answer of the given question. The answer of the converted question must be YES. Do NOT omit the condition in the given question like ’If ...’ or ’Suppose ...’. You MUST include the entire contents of the given question to the converted question.
@@
Table 12: Examples of original datasets and rephrased AR-LSAT.
@@
Dataset Model Det-T Det-F Non-Det
High Low High Low High Low
MusiQue LLaMA3 8B 0.21 0.182 0.141 0.261 0.04 0.166
Mistral 7B 0.159 0.103 0.218 0.27 0.095 0.155
Gemma 7B 0.009 0.022 0.546 0.42 0 0.003
Qwen2 7B 0.25 0.174 0.139 0.327 0.007 0.104
GSM8K LLaMA3 8B 0.209 0.092 0.286 0.414 0 0
Mistral 7B 0.076 0.03 0.407 0.486 0.001 0
Gemma 7B 0.097 0.065 0.378 0.459 0 0
Qwen2 7B 0.194 0.121 0.278 0.407 0 0
AR-LSAT LLaMA3 8B 0.057 0.06 0.303 0.312 0.158 0.11
Mistral 7B 0.058 0.078 0.385 0.437 0.002 0.04
Gemma 7B 0.011 0.025 0.49 0.474 0 0
Qwen2 7B 0.068 0.108 0.306 0.413 0.092 0.013
@@
MusiQue LLaMA3 8B 0.364 0.39 0.268 0.326 0.136 0.197
Mistral 7B 0.432 0.449 0.265 0.28 0.203 0.163
Gemma 7B 0.185 0 0.029 0 0 0
Qwen2 7B 0.122 0.15 0.074 0.093 0 0.032
GSM8K LLaMA3 8B 0.662 0.694 0.629 0.715 N/A N/A
Mistral 7B 0.817 0.667 0.746 0.638 1 N/A
Gemma 7B 0.14 N/A 0.254 N/A N/A N/A
Qwen2 7B 0.315 0.586 0.389 0.386 N/A N/A
AR-LSAT LLaMA3 8B 0.5 0.432 0.298 0.33 0.228 0.281
Mistral 7B 0.585 0.189 0.289 0.195 0 0
Gemma 7B 0.143 0.158 0.146 0.131 N/A N/A
Qwen2 7B 0.22 0.098 0.079 0.095 0.062 0.111
Table 15: Response shift ratio by NASA on false negative samples (FN -> TP ratio) in the rephrased datasets. High denotes high-confident and Low denotes low-confident.
@@
MusiQue LLaMA3 8B 0.057 0.06 0.144 0.112 0 0.06
Mistral 7B 0.073 0.075 0.097 0.08 0.119 0.032
Gemma 7B 0 N/A 0.007 N/A 0 N/A
Qwen2 7B 0.009 0.018 0.015 0.026 0 N/A
GSM8K LLaMA3 8B 0.587 0.632 0.516 0.606 N/A N/A
Mistral 7B 0.784 0.65 0.736 0.652 N/A N/A
Gemma 7B 0.271 N/A 0.188 N/A N/A N/A
Qwen2 7B 0.213 0.208 0.234 0.303 N/A N/A
AR-LSAT LLaMA3 8B 0.229 0.349 0.306 0.247 0.176 0.21
Mistral 7B 0.208 0.204 0.248 0.162 0 0.174
Gemma 7B 0.133 0.083 0.152 0.134 N/A N/A
Qwen2 7B 0.109 0.041 0.13 0.095 0 0
Table 16: Response shift ratio by NASA on true negative samples (TN -> FP ratio) in the rephrased datasets. High denotes high-confident and Low denotes low-confident.

### v2

@@ 6.2 Enhanced Model Calibration
LLaMA 0.594 / 0.596 0.329 / 0.325 0.235 / 0.227
Mistral 0.415 / 0.408 0.126 / 0.133 0.174 / 0.174
Gemma 0.139 / 0.137 0.257 / 0.251 0.125 / 0.131
Qwen 0.541 / 0.541 0.478 / 0.481 0.220 / 0.221
Table 6: Expected calibration error of baseline and ours.
Dataset Mistral + NASA
@@
6.3 Analysis in a Few-shot Scenario
We compare the F1 scores of the baselines and our models in a 4-shot setting on StrategyQA, rephrased GSM8K / AR-LSAT. As shown in Table 7, we first observe that the F1 score gap between the baselines and our models decreases compared to the zero-shot scenario, as reported in Tables 3 and 4. Given that few-shot prompting has been reported to improve model calibration Kadavath et al. (2022), these results suggest that the inclusion of few-shot examples enhances model calibration, thereby partially reducing negative bias. Nevertheless, our model consistently outperforms the baselines in terms of F1 score across all cases. This indicates that NASA remains effective in mitigating negative bias in few-shot scenarios.
7 Discussion
Through our experiments, we identify two key characteristics of negative bias.
Model NAS may not negatively correlate with F1 score during fine-tuning As shown in Table 19, we observe instances where freezing certain parameters in the attention module yields higher F1 scores despite an increase in NAS. These results suggest that a phase during fine-tuning exists where the negative correlation between the F1 score and NAS is disrupted. To mitigate performance degradation, we introduce thresholds such as early stopping and halting within our framework.
Table 7: Precision, recall, and F1 scores in a 4-shot scenario (baseline / NASA).
Model Dataset Precision Recall F1
LLaMA StrategyQA 0.899 / 0.872 0.812 / 0.851 0.853 / 0.862
GSM8K 0.543 / 0.532 0.720 / 0.870 0.619 / 0.661
AR-LSAT 0.562 / 0.557 0.431 / 0.590 0.488 / 0.573
Mistral StrategyQA 0.877 / 0.832 0.824 / 0.889 0.850 / 0.859
GSM8K 0.548 / 0.522 0.763 / 0.885 0.638 / 0.656
AR-LSAT 0.552 / 0.532 0.589 / 0.720 0.570 / 0.612
Gemma StrategyQA 0.903 / 0.897 0.654 / 0.659 0.758 / 0.760
GSM8K 0.588 / 0.590 0.876 / 0.894 0.704 / 0.711
AR-LSAT 0.550 / 0.535 0.366 / 0.405 0.439 / 0.461
Qwen StrategyQA 0.901 / 0.894 0.770 / 0.805 0.830 / 0.847
GSM8K 0.596 / 0.569 0.765 / 0.827 0.670 / 0.674
AR-LSAT 0.535 / 0.546 0.194 / 0.303 0.285 / 0.390
Accurate identification of negative attention heads is crucial, and NAS facilitates this process Our ablation study in Table 20 indicates that fine-tuning randomly selected attention heads leads to suboptimal results. In some cases, we observe abnormal increases or decreases in model NAS, implying the importance of fine-tuning problematic attention heads rather than unrelated parameters. Leveraging NAS and the probing set, we successfully detect negative attention heads, leading to consistent improvements in the F1 score.
@@ 8 Conclusion
We identify a critical issue where large language models exhibit negative bias in binary decision tasks requiring complex reasoning. To address this issue, we propose a negative attention score and employ it to discover query-agnostic negative heads. By performing parameter-efficient tuning on these heads, we introduce the NASA method, which effectively mitigates the bias problem. Our method not only enhances the performance of the model but also serves as a useful analytical framework from the perspective of interpretability.
@@
Acknowledgements
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2022R1A3B1077720, No. 2022R1A5A708390811), Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)], the BK21 FOUR program of the Education and Research Program for Future ICT Pioneers, Seoul National University, Samsung SDS Co., Ltd, and a grant from the Yang Young Foundation.
@@ Appendix A Data Statistics
Table 8: Data statistics of datasets utilized in our work.
@@
Appendix B A Statistical Analysis of the Relationship between Negative Confidence and NAS
Table 9: Results of multiple regression alongside a correlation for multiple comparisons. $r^{2}$ denotes the effect size.
Model Coefficient Corrected $p$-value $r^{2}$
[…]
To analyze the impact of negative confidence on model NAS, we conduct a multiple regression analysis by including positive confidence as a variable, as shown in Table 9. In all cases except for LLaMA, where the corrected $p$-value is high, negative confidence is observed to have a statistically significant positive correlation with model NAS. Meanwhile, considering the effect sizes, future work could investigate additional factors influencing model NAS.
Appendix C Instructions for NASA
@@
Appendix D Details of Binary Decision Data Construction
@@
D.1 GPT-4 Prompts for Negative Example Construction
@@
Appendix E Details of Parametric Sample Selection
As mentioned in Section 4.1, we construct the probing and fine-tuning sets using the HotpotQA dataset, which is a short-answer QA dataset. From the HotpotQA dataset, we use samples that can be correctly answered solely with parametric knowledge, without any context, as our probing samples. Specifically, we suppose that the probing samples must meet the condition: the question should be answerable using the model’s parametric knowledge. To determine whether the question in the sample meets the condition, we construct the prompt as follows:
This prompt is designed to identify attention heads that attend to the negative token in the instruction even when a positive answer is expected. We assume that if attention heads attend to the negative token when the model knows the correct answer and the target answer is “Yes”, it is generated by a negative bias in the model. This prompt effectively selects samples that belong to the model’s parametric knowledge while minimizing the influence of the model’s negative bias. We assume that samples meet the condition if the model returns the correct answers for the prompt, and we use them as the probing set. We utilize GPT-4 to verify whether the prediction matches the label.
@@ Appendix E Details of Parametric Sample Selection
Appendix F Algorithm of Incremental Head Tuning
Algorithm 1 shows the details of head-wise incremental tuning of NASA.
Appendix G Case Study for Negative Responses
@@
Algorithm 1 Head-wise Incremental Tuning
Input: Fine-tuning set $D$, validation set $V$, model $\psi$, single head NAS threshold $\rho$, list of negative attention heads to be fine-tuned $P$, early halting threshold $\tau$, maximum training epoch $T$, set of all layer and head indices $L$ and $H$, respectively.
Output: NASA-tuned model.
1:for each $(l,h)\in P$ do
2:     $\psi_{init}=\psi$
3:     $\alpha_{init}=\text{NAS}(V,l,h)$
4:     $\beta_{init}=\text{NAS}(V,L,H)$
5:     $\alpha=\beta=\infty$
6:     for epoch $=1$ to $T$ do
7:         Update $\psi$ by fine-tuning $h$-th
8:attention head in the $l$-th layer using $D$.
9:         $\alpha^{\prime}=\text{NAS}(V,l,h)$
10:         $\beta^{\prime}=\text{NAS}(V,L,H)$
11:         if $\alpha^{\prime}>\alpha$ or $\alpha^{\prime}<\rho$ or $\beta^{\prime}>\beta$ then
12:              break $\triangleright$ Early stopping
13:         $\alpha\leftarrow\alpha^{\prime},\beta\leftarrow\beta^{\prime}$
14:     if $\alpha^{\prime}>\alpha_{init}$ or $\beta^{\prime}>\beta_{init}$ then
15:         $\psi\leftarrow\psi_{init}$triangleright Update cancellation
16:     if $\beta^{\prime}<\tau$ then
17:         break $\triangleright$ Early halting
18:return $\psi$
@@
We apply this classification to cases where the model gives a negative response in a binary decision task. Negative responses can be classified as true negative (TN) or false negative (FN), with our focus in this discussion on FN. In this, an FN response classified as high-confident Det-T is the case where the model knows the answer is correct but says “No”, while an FN response classified as Non-Det is the case where the model doesn’t know the correct answer and says “No”. For high-confident Det-F responses, the model does not indicate unanswerability but rather holds incorrect knowledge. This issue falls outside the scope of our methodology, which focuses on handling negative bias in binary decision tasks rather than general knowledge-intensive tasks.
@@
Although it is clear that the case where the model knows the answer is correct but says “No” is the most problematic, we argue that in a binary decision task, the case where the model doesn’t know the correct answer and says “No” can be related to the issue of negative bias. In a binary decision task where the model must answer “Yes” or “No”, instances where the model is uncertain about the correct answer should result in an equal frequency of positive (yes) and negative (no) responses. However, our analysis indicates that existing models tend to output negative responses more confidently and frequently (as shown in Figures 1 and 2 in the paper). This suggests that when the model is uncertain, it predominantly outputs negative responses, which negatively impacts the trustworthiness of these negative responses. Therefore, for models exhibiting such negative bias, Non-Det and low-confident responses should also be partially adjusted.
Table 16 shows the compositions of FN responses from the original models according to the classifications mentioned. It is evident that across all types of reasoning tasks examined in our work, including multi-hop, mathematical, and logical reasoning, FN responses encompass two distinct cases: (A) instances where the model recognizes the correct answer but incorrectly responds with “No”, and (B) instances where the model lacks knowledge of the correct answer and responds with “No”.
Meanwhile, for Gemma, we observe that the model tends to generate overly cautious responses when performing multi-hop QA. Specifically, instead of selecting the “unanswerable” option for multiple samples, Gemma rephrases its responses in an alternative long-form manner. As a result, instances that should have originally been classified as Non-Det are instead categorized as Det-F. This should be taken into account when interpreting the results in Table 16.
Appendix H Analysis of Negative Responses Shifting in NASA
@@
In this section, we analyze the ratios of samples which are originally responded to as negative by the original model and subsequently changed to positive by the NASA model. We categorize the ratio cases by each classification group explained in section G.
@@
Table 17 illustrates the ratios of samples shifted from FN to TP. This demonstrates that the Det-T corresponding to the case where the model knows the answer is correct but says “No” generally has the highest rate of change. This finding indicates that our methodology does not indiscriminately add positive bias to the model but rather enhances the model’s binary decision reasoning ability effectively.
@@
Table 18 illustrates the ratios of samples shifted from TN to FP. As discussed in the section on main results, some samples that are incorrectly categorized as TN due to negative bias might shift to FP (Det-F), and this case does not correspond to the model degradation. Actually, among the TN samples, those classified as Det-T with high confidence exhibit a lower shift ratio compared to their Det-F counterparts. This indicates that the model tends to maintain predictions for samples that it can accurately classify. In conclusion, it can be interpreted that NASA has improved reasoning capability in binary decision tasks while maintaining general reasoning capability.
Appendix I Generalization to Universal Binary Decision
We analyze the generalization ability for binary decisions by replacing the positive and negative vocabulary used in our fine-tuning set. Our fine-tuning set consists of “Yes” and “No” in the instruction, and we test with the tokens “True” and “False”, as well as “Correct” and “Wrong”.
Type Dataset LLaMA + NASA
True/False StrategyQA 0.855 / 0.828 0.866 / 0.853
MATH 0.559 / 0.452 0.574 / 0.583
AR-LSAT 0.512 / 0.306 0.506 / 0.405
Correct/Wrong StrategyQA 0.852 / 0.838 0.846 / 0.840
MATH 0.545 / 0.660 0.544 / 0.675
AR-LSAT 0.523 / 0.540 0.522 / 0.590
Table 10 presents the results of experiments with LLaMA on three datasets transformed for the binary decision task. We observe that models fine-tuned with NASA achieve semantic-level generalization without overfitting to the specific vocabulary.
Appendix J Transferability across Various Instructions
It. Type Dataset LLaMA + NASA
A MuSiQue 0.686/0.641 0.710/0.700
GSM8K 0.534/0.472 0.543/0.575
AR-LSAT 0.508/0.306 0.507/0.370
B MuSiQue 0.708/0.692 0.736/0.756
GSM8K 0.536/0.611 0.531/0.665
AR-LSAT 0.512/0.382 0.518/0.500
We fix the instruction for binary decision-making throughout the application of our method. In this section, we evaluate the robustness of our method when presented with instructions in the inference stage that are different forms of content. We assess the binary decision performance using the LLaMA with and without NASA on benchmarks: MuSiQue, GSM8K, and AR-LSAT using paraphrased instruction types A and B. Note that the datasets are transformed for the binary decision task. Details of instruction types can be found in Appendix C. As shown in Table 11, models fine-tuned with NASA still demonstrate superior performance compared to the baseline in both prompt types. This supports our claim of instruction robustness for our method. Fig 11 in Appendix C shows the content of instructions.
Appendix K Ablation Study
@@ Appendix I Generalization to Universal Binary Decision
K.1 Freezing Key Projection Weight
@@
The experimental results of NASA and freezing key projection weights are shown in Table 19. In general, both the F1 score and NAS are relatively better for NASA, while the trend in accuracy is not distinctly observable. An exceptional case is Qwen at AR-LSAT, where NASA’s accuracy and F1 score are significantly lower than those of its counterpart.
K.2 Random Attention Head Tuning
In another ablation study, we experiment by randomly selecting attention heads to be fine-tuned. As shown in Table 20, NAS typically outperforms most others in terms of F1 score and expected calibration error. Notably, during random attention head tuning, model NAS largely diverges in the case of LLaMA and Mistral. These results imply that carefully selecting attention heads is critical and our framework effectively detects and addresses negatively biased attention heads.
Appendix L Confidence Histogram Analysis
To analyze the confidence of binary decisions between the baseline and our model, we plot confidence histograms for the LLaMA model, similar to Fig 2. In Fig 5, we observe that NASA influences the prediction confidence of LLMs in binary decision tasks. In the baseline model, except for StrategyQA, we frequently observe the return of negative responses with high confidence. On the other hand, after incorporating NASA, the model demonstrates a reduction in the frequency and confidence gap between positive and negative responses.
@@
Appendix M Samples of Attention Heads
Fig 6 presents the real examples of attention weights of heads in the LLaMA model for a GSM8K rephrased sample. Fig 7 presents the real examples of attention weights of heads in the LLaMA model with NASA for the same GSM8K rephrased sample.
Appendix N Computation
@@
Appendix O Licenses
StrategyQA, MuSiQue, MetaMATH, and AR-LSAT are under the license of MIT license, CC-BY-4.0 license, MIT License, and MIT license, respectively. LLaMA3-8B-Instruct, Mistral-7B-Instruct-v0.3, Gemma-1.1-7b-it, and Qwen2-7B-Instruct, and GPT-4 are under the license of META LLAMA 3 COMMUNITY LICENSE AGREEMENT, Apache License 2.0, Gemma, Apache License 2.0, and OpenAI, respectively.
Appendix P Usage of AI Writing Assistance
@@
Table 12: Instructions used NASA process in our work and Table 11.
@@
Convert the given question to the binary Yes-No question based on the context about the given question and the answer of the given question. The answer of the converted question must be YES. Do NOT omit the condition in the given question like ‘If ...’ or ‘Suppose ...’. You MUST include the entire contents of the given question to the converted question.
@@
Table 14: Examples of original datasets and rephrased AR-LSAT. Examples are sampled from StrategyQA (Geva et al., 2021), MuSiQue (Trivedi et al., 2022), GSM8K-Rephrased and MATH-Rephrased (Yu et al., 2023), and AR-LSAT (Zhong et al., 2021).
@@
MuSiQue LLaMA 0.21 0.182 0.141 0.261 0.04 0.166
Mistral 0.159 0.103 0.218 0.27 0.095 0.155
Gemma 0.009 0.022 0.546 0.42 0 0.003
Qwen 0.25 0.174 0.139 0.327 0.007 0.104
GSM8K LLaMA 0.209 0.092 0.286 0.414 0 0
Mistral 0.076 0.03 0.407 0.486 0.001 0
Gemma 0.097 0.065 0.378 0.459 0 0
Qwen 0.194 0.121 0.278 0.407 0 0
AR-LSAT LLaMA 0.057 0.06 0.303 0.312 0.158 0.11
Mistral 0.058 0.078 0.385 0.437 0.002 0.04
Gemma 0.011 0.025 0.49 0.474 0 0
Qwen 0.068 0.108 0.306 0.413 0.092 0.013
Table 17: Response shift ratio by NASA on false negative samples (FN $\rightarrow$ TP ratio) in the rephrased datasets. High denotes high-confident and Low denotes low-confident.
@@
MuSiQue LLaMA 0.364 0.39 0.268 0.326 0.136 0.197
Mistral 0.432 0.449 0.265 0.28 0.203 0.163
Gemma 0.185 0 0.029 0 0 0
Qwen 0.122 0.15 0.074 0.093 0 0.032
GSM8K LLaMA 0.662 0.694 0.629 0.715 N/A N/A
Mistral 0.817 0.667 0.746 0.638 1 N/A
Gemma 0.14 N/A 0.254 N/A N/A N/A
Qwen 0.315 0.586 0.389 0.386 N/A N/A
AR-LSAT LLaMA 0.5 0.432 0.298 0.33 0.228 0.281
Mistral 0.585 0.189 0.289 0.195 0 0
Gemma 0.143 0.158 0.146 0.131 N/A N/A
Qwen 0.22 0.098 0.079 0.095 0.062 0.111
Table 18: Response shift ratio by NASA on true negative samples (TN $\rightarrow$ FP ratio) in the rephrased datasets. High denotes high-confident and Low denotes low-confident.
Dataset Model Det-T Det-F Non-Det
High Low High Low High Low
MuSiQue LLaMA 0.057 0.06 0.144 0.112 0 0.06
Mistral 0.073 0.075 0.097 0.08 0.119 0.032
Gemma 0 N/A 0.007 N/A 0 N/A
Qwen 0.009 0.018 0.015 0.026 0 N/A
GSM8K LLaMA 0.587 0.632 0.516 0.606 N/A N/A
Mistral 0.784 0.65 0.736 0.652 N/A N/A
Gemma 0.271 N/A 0.188 N/A N/A N/A
Qwen 0.213 0.208 0.234 0.303 N/A N/A
AR-LSAT LLaMA 0.229 0.349 0.306 0.247 0.176 0.21
Mistral 0.208 0.204 0.248 0.162 0 0.174
Gemma 0.133 0.083 0.152 0.134 N/A N/A
Qwen 0.109 0.041 0.13 0.095 0 0